OpenStudy (anonymous):

compute lim x->0 (x^2(sin(4/x)))

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} x^2 \sin(4/x)\]

OpenStudy (anonymous):

Show steps please

OpenStudy (freckles):

how about squeeze theorem

OpenStudy (aum):

The sine function is confined to [-1, 1]. Therefore,\[ -\lim_{x \rightarrow 0} x^2 ~~\le ~~ \lim_{x \rightarrow 0} x^2 \sin(4/x) ~~\le ~~ \lim_{x \rightarrow 0} x^2 \\ -0 ~~\le ~~ \lim_{x \rightarrow 0} x^2 \sin(4/x) ~~\le ~~ +0\\ \lim_{x \rightarrow 0} x^2 \sin(4/x) = 0 \]

OpenStudy (freckles):

pretty

OpenStudy (aum):

thanks.

OpenStudy (anonymous):

is that how I show that the limit equals 0 on an exam?

OpenStudy (zarkon):

it depends on how picky your instructor is...look at the following for an example http://openstudy.com/users/zarkon#/updates/54825bcfe4b0edc680fcf779 and look at what loser66 got when he did a squeeze theorem problem

OpenStudy (aum):

I can add two more steps to my earlier reply just for clarification. The sine function is confined to [-1, 1]. Therefore,\[ x^2*(-1) ~~\le ~~ x^2 * \sin(4/x) ~~\le ~~ x^2 * (1) \\ -x^2 ~~\le ~~ x^2 * \sin(4/x) ~~\le ~~ x^2\\ \text{ } \\ \lim_{x \rightarrow 0} -x^2 ~~\le ~~ \lim_{x \rightarrow 0} x^2 \sin(4/x) ~~\le ~~ \lim_{x \rightarrow 0} x^2 \\ -0 ~~\le ~~ \lim_{x \rightarrow 0} x^2 \sin(4/x) ~~\le ~~ +0\\ \lim_{x \rightarrow 0} x^2 \sin(4/x) = 0 \]